OpenStudy (anonymous):

Let f(x) = 2 x + 4 and g(x) = 3 x^2 + 5 x. After simplifying, (f+g)(x) =

OpenStudy (anonymous):

Welcome to OpenStudy! @statc1

OpenStudy (anonymous):

Do you know that \(\sf\Large (f+g)(x) = f(x)+g(x)\)

OpenStudy (anonymous):

So substitute the values you were given and tell me what you get.

OpenStudy (anonymous):

f(x)+g(x)=2x+4+3x^2+5x=3x^2+7x+4 =3x^2+3x+4x+4 =3x(x+1)+4(x+1)=(x+1)(3x+4) answers